# End-to-End Autonomous Driving through V2X Cooperation

### v2

@@ 3.6 Cross-View Data Fusion (Occupancy Fusion)
3.7 Planning Output
Based on the fused agent queries, lane queries, and occupancy features, we reused the implementations in UniAD [16] to generate final planning output. MotionFormer is used to generate a set of $N_{a}$ motion queries with a prediction horizon of $t_{pred}$. These queries are created by capturing the interactions among agents, lanes, and goals. Notably, these agent queries encompass the ego-vehicle query, thereby enabling MotionFormer to generate ego-vehicle queries with multimodal intentions. The BEV occupied probability map $\hat{P}$ is utilized to create a binary occupancy map $\hat{O}$. In the planning phase, the ego-vehicle query obtained from MotionFormer is combined with command embeddings to shape a "plan query". These commands comprise turning left, turning right, and moving forward. This plan query, along with the BEV feature, is input into the decoder to produce future waypoints. The ultimate planning trajectory is derived by minimizing a cost function, ensuring the avoidance of collisions with occupied grids $\hat{O}$.
3.8 UniV2X Training
In our work, we train UniV2X with imitation learning. Specifically, to ensure comprehensive and stable training, we adopt a four-stage training strategy. In the first stage, we pre-train the infrastructure system, including tracking, online mapping, and occupancy prediction tasks, with the annotations from infrastructure view as ground truth. In the second stage, we focus on pre-training perception tasks within the ego-vehicle system. This stage encompasses tasks like tracking and online mapping, employing the vehicle-view annotations as ground truth. In the third stage, all fusion modules and all modules in ego-vehicle systems, as well as all tasks, are trained. Cooperative-view annotations are used as the ground truth for this stage, with the following combined losses:
$L=L_{track}+L_{map}+L_{motion}+L_{occ}+L_{plan}.$ (5)
Furthermore, we use self-supervised learning to train QueryFlowNet and OccFlowNet. This involves constructing infrastructure frame pairs and employing a similarity loss for training, similar to the strategy described in  [62].
4 Experiments
In this section, we implement UniV2X, alongside reproducing various perception, online mapping, and end-to-end methods on DAIR-V2X [60, 63]. Experiment results showcase the effectiveness, transmission-friendliness, and reliability of our design. More details, ablation studies and visualizations are provided in the Appendix. Furthermore, we conduct UniV2X on more V2X datasets such as V2X-Sim [25], and present the experiment results in the Appendix.
Table 3: Planning Evaluation Results. We do not report the results at 0.5s and 1.5s because most of the collision rate is zero. UniV2X significantly surpasses all reproduced end-to-end baselines in collision rate, while requiring much less transmission cost.
@@ 4 Experiments
No Fusion 2.58 3.37 4.36 3.44 0.15 1.04 1.48 0.89 0
Vanilla 2.21 3.31 4.46 3.33 0.15 0.89 2.67 1.24 8.19×$\times 10^{7}$
V2VNet[47] 2.31 3.29 4.31 3.30 0.00 1.03 1.47 0.83 8.19×$\times 10^{7}$
CooperNaut [9] 3.83 5.26 6.69 5.26 0.59 1.92 1.63 1.38 8.19×$\times 10^{7}$
UniV2X (Ours) 2.60 3.34 4.36 3.43 0.00 0.74 0.74 0.49 8.09×$\times 10^{5}$
4.1 Experiment Settings
DAIR-V2X Dataset comprises approximately 100 scenes captured at 28 complex traffic intersections, recorded using both infrastructure and vehicle sensors. Each scene has a duration ranging from 10 to 25 seconds, capturing data at a rate of 10 Hz, and is equipped with a high-definition (HD) map. This dataset provides a diverse range of driving behaviors, including actions such as moving forward, turning left, and turning right. To align with nuScenes [3], we categorize object classes into four categories (car, bicycle, pedestrian, traffic_cone).
UniV2X Implementation Details. We establish the interest range of the ego vehicle as [-50, 50, -50, 50] meters. The ego-vehicle BEV range is defined within a rectangular area spanning [-50, 50, -50, 50] meters, with each grid measuring 0.25m by 0.25m. The infrastructure BEV range is set as [0, 100, -50, 50] meters, accounting for the camera’s forward sensing range and facilitating more effective utilization of infrastructure data. The experiments are conducted utilizing 8 NVIDIA A100 GPUs. More implementation details can be found in the appendix.
Baseline Settings. No Fusion only utilizes ego-vehicle images as sensor data input, without any infrastructure data input. In Vanilla approach, we employ a simple CNN to fuse infrastructure and ego-vehicle BEV features. The fused BEV feature is reshaped into one dimension and subsequently fed into a Multi-Layer Perceptron (MLP) to generate the planning path. In V2VNet [47] approach, we use a CNN to fuse two-side BEV features into a new ego-vehicle BEV feature, and send this new feature into UniAD [16]. In CooperNaut [9], we also employ a CNN to fuse two-side BEV features, with only one frame input each time. Given the significant role of ego status, such as ego-vehicle velocity, in open-loop end-to-end autonomous driving, as illustrated in [29], we remove the ego-vehicle velocity embedding in all baseline settings for a fair comparison. Additionally, we explore the role of ego-vehicle velocity for UniV2X in the appendix.
4.2 Experiment Results on DAIR-V2X
4.2.1 Planning Results.
We report the planning results in Table 3. Compared with No Fusion, UniV2X achieves a 45% decreasing in average collision rate in Table 3. In particular, as the planning time increases, the performance improvement becomes more obvious. This result effectively illustrates that the use of infrastructure information can improve autonomous driving performance, especially for low-cost monocular solutions. Compared with Vanilla, V2VNet methods, and CooperNaut, UniV2X behaves much better in average collision rate (0.49 vs 1.24, 0.49 vs 0.83, 0.49 vs 1.38). Note that we ignore the evaluation result comparisons in L2 Error since most methods achieve similar L2 Error results. Furthermore, UniV2X requires much less transmission cost than Vanilla as well as V2VNet and CooperNaut solutions ($8.09\times 10^{5}$ vs $8.19\times 10^{7}$). So that our UniV2X requires much less transmission cost and is more transmission-friendly.
4.2.2 Agent Perception Results.
We employ various fusion strategies on DAIR-V2X, including No Fusion (no fusing infrastructure data), Early Fusion (fusing raw infrastructure BEV feature), and Late Fusion (fusing infrastructure detection results with Hungarian method[23]). Additionally, we reproduce current SOTA cooperative perception methods on DAIR-V2X, namely V2X-ViT [52], Where2comm [17], and DiscoNet [26]. For a fair comparison, we standardize inputs (image-only) and evaluation settings. All methods, except for CoCa3D [18] based on depth estimation, are re-implemented using BEVFormer [27].
@@ 4.2 Experiment Results on DAIR-V2X
[…]
@@
Late Fusion+AB3DMOT [49] 0.196 0.263 6.60×$\times 10^{2}$
CoCa3D [18] 0.226 - 4.63×$\times 10^{6}$
V2X-ViT [52] 0.268 0.287 2.56×$\times 10^{6}$
Where2comm [17] 0.162 0.106 5.40×$\times 10^{5}$
DiscoNet [26] 0.216 0.203 1.60×$\times 10^{5}$
V2X-ViT [52]+Where2comm [17] 0.178 0.071 7.22×$\times 10^{4}$
@@
We present the evaluation results (car class) for detection and tracking in Table 4. (1) UniV2X demonstrates a notable enhancement of +7.6 and +3.0 in AMOTA(%) compared to No Fusion and Early Fusion. (2) UniV2X outperforms CoCa3D, Where2comm, and DiscoNet at similar or less transmission cost. (3) While UniV2X follows an end-to-end paradigm, achieving inferior tracking performance compared to tracking-by-detection methods with complex association, such as Late Fusion+AB3DMOT (0.239 vs 0.263 at AMOTA), it significantly outperforms this tracking-by-detection solution in detection (0.295 vs 0.196). It is important to note that this tracking-by-detection solution is not suitable for end-to-end autonomous driving. (4) V2X-ViT exhibits better performance than UniV2X at AMOTA (0.287 vs 0.239), but it requires much more transmission cost (2.56×$\times 10^{6}$ vs 6.94×$\times 10^{4}$). When we further compress the V2X-ViT transmission to a level similar to UniV2X with Where2comm, there is a significant performance drop (from 0.287 to 0.071 at AMOTA). These outcomes underscore the capability of our infrastructure agent queries and agent fusion module in enhancing agent perception ability with light transmission cost.
4.2.3 Online Mapping Results.
We implement No Fusion, Early Fusion, and CoBEVT[51] for online mapping on DAIR-V2X. All methods are re-implemented using BEVFormer [27]. The mapping performance is reported with Segmentation Intersection over Union (IoU) (%) as the evaluation metric in Table 5. UniV2X demonstrates notable improvements in lane perception and crossing perception compared No Fusion, Early Fusion and CoBEVT, respectively. Moreover, compared with Early Fusion and CoBEVT, UniV2X requires less than 1/10th of the transmission cost. These results indicate that infrastructure lane queries and cross-view lane fusion are effective in enhancing online mapping ability.
@@
CoBEVT [51] 15.6 16.4 2.56×$\times 10^{6}$
UniV2X (Ours) 17.8 19.8 1.47×$\times 10^{5}$
@@
4.2.4 Occupancy Prediction Results.
Concerning the evaluation of occupancy prediction, as depicted in Table 6, UniV2X exhibits notably superior performance compared to No Fusion in both near and far regions. Particularly, UniV2X achieves +5.7 and +13.4 improvement in IoU-n (%) and IoU-f (%) respectively. Here, "IoU-n" and "IoU-f" denote evaluation ranges of 30×30m and 50×50m, respectively. These results underscore the effectiveness of our sparse-dense hybrid transmitted data in significantly enhancing occupancy prediction.
4.3 Ablation Study
4.3.1 Roles of Each Fusion Module.
In this part, we investigate the contributions of each fusion module within UniV2X to the enhancement of final planning performance. In evaluating the trained UniV2X, we only fuse agent queries, lane queries, and occupancy probability maps, respectively. We assess each individual fusion module and present the results in Table 7.
As evident in Table 7, AgentFusion and OccFusion exhibit an average collision rate decrease of 0.15% and 0.30%, respectively, while LaneFusion shows a marginal decrease in the average collision rate. Notably, when AgentFusion and OccFusion are combined, there is a more pronounced decrease in the collision rate. These results emphasize the effectiveness of AgentFusion and OccFusion in enhancing collision avoidance performance. However, it is crucial to note that the lack of a substantial decrease in the collision rate with LaneFusion does not imply its ineffectiveness for final planning. Metrics such as the lane violation rate [64] are essential for measuring planning output, reflecting the map enhancement for fewer collisions with lane lines and better adherence to traffic rules. In our future work, we plan to incorporate additional evaluation metrics for a more comprehensive assessment of planning performance.
Table 7: Planning Evaluation Results of UniV2X with Each Fusion Module. AgentFusion and OccFusion both play a role in reducing the collision rate.
Method L2 Error ($m$$\downarrow$ Col. Rate(%$\downarrow$
@@ 4.3 Ablation Study
No Fusion 2.58 3.37 4.36 3.44 0.15 1.04 1.48 0.89
AgentFusion 2.55 3.35 4.34 3.41 0.30 1.04 0.89 0.74 (-0.15)
LaneFusion 2.58 3.37 4.36 3.44 0.15 1.04 1.48 0.89 (-0.0)
OccFusion 2.63 3.37 4.38 3.46 0.00 0.59 1.19 0.59 (-0.30)
UniV2X 2.60 3.34 4.36 3.43 0.00 0.74 0.74 0.49 (-0.40)
4.3.2 Reliability in Different Communication Conditions.
Figure 4: Reliability on Data Corruption.
When assessing UniV2X, our initial step involves randomly discarding 10%, 30%, 50%, 70%, and 100% of infrastructure agent queries during transmission to simulate data corruption. Following this, the retained queries are utilized for cross-view query transmission and interaction, and the performance of agent perception, encompassing object detection and tracking, is evaluated accordingly. The evaluation results depicted in Figure 4 unveil a gradual degradation in agent perception as the data corruption ratio increases. When the data corruption ratio reaches 100%, meaning only ego-vehicle sensor data can be used, the performance of UniV2X becomes comparable to that of the No Fusion model. This decline in performance is anticipated, as data corruption diminishes the complementary information crucial for ego-vehicle autonomous driving. Moreover, it is noteworthy that even in the absence of certain data due to corruption, UniV2X can maintain a basic level of performance comparable to the No Fusion model. This underscores the reliability of our transmission and cross-view data fusion mechanism.
5 Conclusion
Limitations and Future Work. The framework involves multiple modules and different agent perspectives, resulting in a high degree of complexity. As a result, several interaction fusion modules within the framework remain in preliminary stages. Further refinement is essential for optimizing the internal design of the subsequent framework. Moreover, the current planning evaluations only adopt L2 Errors and Collision Rate metrics. In the future, more comprehensive evaluation metrics will be adopted to measure the planning performance. In this work, we only consider open-loop evaluation for end-to-end autonomous driving. We will conduct more closed-loop experiments to evaluate our UniV2X.
Appendix 0.A Experiment Results on More Dataset
In this section, we apply UniV2X to the V2X-Sim dataset [25] to showcase its effectiveness across a broader spectrum of datasets and V2X scenarios.
0.A.0.1 V2X-Sim Dataset.
The V2X-Sim dataset [25] is a synthetic dataset for cooperative autonomous driving, developed through the integration of Carla Simulator [11] and SUMO [32] co-simulation. This dataset comprises 100 scenes, and each scene contains synchronized image and point cloud sequences, at a key frame rate of 5Hz, from roadside sensors and sensors mounted on five autonomous driving vehicles. The dataset also provides a diverse array of meticulously annotated ground truths, including 3D bounding boxes and tracking IDs.
0.A.0.2 Experiment Settings and Results.
We choose two autonomous vehicles from V2X-Sim and designate one as the ego vehicle to establish V2V scenes. To maintain consistency with the implementation on DAIR-V2X [60], we equip each autonomous vehicle solely with a front camera sensor. For this experiment, we utilize only one-fifth of the dataset to streamline the results. Considering the absence of evaluation metrics reflecting online mapping’s impact on planning, such as lane violation, we exclude the online mapping task from this experiment. We proceed with training and evaluation of UniV2X on V2X-Sim using settings akin to those in DAIR-V2X [60]. The experimental outcomes are detailed in Table 8, Table 10, and Table 10.
Table 8: Planning Evaluation Results on V2X-Sim Dataset [25].
Method L2 Error ($m$$\downarrow$ Col. Rate (%$\downarrow$

### v3

@@ Introduction
Planning Output
With the fused agent queries, lane queries, and occupancy features, we first generate rough future waypoints by reusing the implementations in UniAD (Hu et al. 2023b). MotionFormer is used to generate a set of $N_{a}$ motion queries with a prediction horizon of $t_{pred}$. These queries are created by capturing the interactions among agents, lanes, and goals. Notably, these agent queries encompass the ego-vehicle query, thereby enabling MotionFormer to generate ego-vehicle queries with multimodal intentions. The BEV occupied probability map $\hat{P}$ is utilized to create a binary occupancy map $\hat{O}$. In the planning phase, the ego-vehicle query obtained from MotionFormer is combined with command embeddings to shape a ”plan query”. These commands comprise turning left, turning right, and moving forward. This plan query, along with the BEV feature, is input into the decoder to produce future waypoints.
The final planning trajectory is determined by: 1) adjusting the future waypoints on the roads to ensure adherence to traffic rules and staying within driving areas with the generated lanes and other road elements, and 2) minimizing a cost function to avoid collisions with occupied grids $\hat{O}$.
Experiments
In this section, we implement UniV2X, alongside reproducing various perception, online mapping, and end-to-end methods on DAIR-V2X (Yu et al. 2022, 2023c). More implementation details, ablation studies, visualizations and analysis are provided in the Appendix. We also conduct UniV2X on more V2X datasets such as V2X-Sim (Li et al. 2022a), and present the experiment results in the Appendix.
Table 2: Planning Evaluation Results. We do not report the results at 0.5s and 1.5s because most of the collision rate is zero. Although CooperNaut achieves a lower off-road rate, it has a much larger L2 error compared to other methods. This is because its planning length is relatively conservative, ensuring it is easier to remain within the drivable area over a given period.
Method L2 Error ($m$$\downarrow$ Col. Rate (%$\downarrow$ Off-Road Rate (%$\downarrow$ Transm. Cos (BPS)$\downarrow$
@@
No Fusion 2.58 3.37 4.36 3.44 0.15 1.04 1.48 0.89 0.44 0.59 2.22 1.08 0
Vanilla 2.33 3.69 5.12 3.71 0.59 2.07 3.70 2.12 0.15 1.33 4.74 2.07 8.19×$\times 10^{7}$
BEV Feature Fusion 2.31 3.29 4.31 3.30 0.0 1.04 1.48 0.83 0.44 0.44 1.91 0.93 8.19×$\times 10^{7}$
CooperNaut (Cui et al. 2022) 3.84 5.33 6.87 5.35 0.44 1.33 1.93 1.23 0.15 0.15 1.33 0.54 8.19×$\times 10^{7}$
UniV2X (Ours) 2.59 3.35 4.49 3.48 0.0 0.44 0.59 0.34 0.74 0.74 1.19 0.89 8.09×$\times 10^{5}$
Experiment Settings
DAIR-V2X Dataset
comprises approximately 100 scenes captured at 28 complex traffic intersections, recorded using both infrastructure and vehicle sensors. Each scene has a duration ranging from 10 to 25 seconds, capturing data at a rate of 10 Hz, and is equipped with a high-definition (HD) map. This dataset provides a diverse range of driving behaviors, including actions such as moving forward, turning left, and turning right. To align with nuScenes (Caesar et al. 2020), we categorize object classes into four categories (car, bicycle, pedestrian, traffic_cone).
Implementation.
We establish the interest range of the ego vehicle as [-50, 50, -50, 50] meters. The ego-vehicle BEV range shares the same area spanning [-50, 50, -50, 50] meters, with each grid measuring 0.25m by 0.25m. The infrastructure BEV range is set as [0, 100, -50, 50] meters, accounting for the camera’s forward sensing range and facilitating more effective utilization of infrastructure data. The experiments are conducted utilizing 8 NVIDIA A100 GPUs. More implementation details are provided in the appendix.
Baseline Settings.
No Fusion only utilizes ego-vehicle images as sensor data input, without any infrastructure data input. In Vanilla approach, we employ a simple CNN to fuse infrastructure and ego-vehicle BEV features. The fused BEV feature is reshaped into one dimension and subsequently fed into a Multi-Layer Perceptron (MLP) to generate the planning path. In BEV Feature Fusion, we use a CNN to fuse two-side BEV features into a new ego-vehicle BEV feature, and send this new feature into UniAD (Hu et al. 2023b). CooperNaut (Cui et al. 2022) originally employs Point Transformer to aggregate cross-view feature by using the sparse characteristics of point clouds. However, the image is a dense representation, and conducting similar sparse operations, as seen with Where2comm (Hu et al. 2022) in Table 3, results in poor performance. Therefore, we directly transmit dense BEV features and use CNN to fuse features from both sides, with only one frame input each time to achieve better comparison. Given the significant role of ego status, such as ego-vehicle velocity, in open-loop end-to-end autonomous driving, as illustrated in (Li et al. 2024), we remove the ego-vehicle velocity embedding in all baseline settings for a fair comparison. Additionally, we explore the role of ego-vehicle velocity for UniV2X in the appendix.
Experiment Results on DAIR-V2X
Planning Results.
@@
Agent Perception Results.
We employ various fusion strategies on DAIR-V2X, including No Fusion, Early Fusion (fusing raw infrastructure BEV feature), and Late Fusion (fusing infrastructure detection results with Hungarian method(Kuhn 2010)). Additionally, we reproduce current SOTA cooperative perception methods on DAIR-V2X, namely V2X-ViT (Xu et al. 2022b), Where2comm (Hu et al. 2022), DiscoNet (Li et al. 2021), and CoAlign (Lu et al. 2023). For a fair comparison, we standardize inputs (image-only) and evaluation settings. All methods, except for CoCa3D (Hu et al. 2023a) based on depth estimation, are re-implemented using BEVFormer (Li et al. 2022b).
@@
Method mAP $\uparrow$ AMOTA $\uparrow$ Trans. Cost $\downarrow$
@@
Late Fusion 0.196 0.263 6.60×$\times 10^{2}$
CoAlign 0.240 0.234 8.19×$\times 10^{7}$
CoCa3D 0.226 - 4.63×$\times 10^{6}$
V2X-ViT 0.268 0.287 2.56×$\times 10^{6}$
Where2comm 0.162 0.106 5.40×$\times 10^{5}$
DiscoNet 0.216 0.203 1.60×$\times 10^{5}$
V2X-ViT+Where2comm 0.178 0.071 7.22×$\times 10^{4}$
@@
We present the evaluation results (car class) for detection and tracking in Table 3. (1) UniV2X demonstrates a notable enhancement of +7.6 and +3.0 in AMOTA(%) compared to No Fusion and Early Fusion. (2) UniV2X outperforms CoCa3D, Where2comm, and DiscoNet at similar or less transmission cost. (3) UniV2X achieves inferior tracking performance compared to tracking-by-detection methods with complex association, such as Late Fusion+AB3DMOT (0.239 vs 0.263 at AMOTA), but it significantly outperforms this tracking-by-detection solution in detection (0.295 vs 0.196). It is important to note that this tracking-by-detection solution is not suitable for end-to-end autonomous driving. (4) V2X-ViT exhibits better performance than UniV2X at AMOTA (0.287 vs 0.239), but it requires much more transmission cost (2.56×$\times 10^{6}$ vs 6.94×$\times 10^{4}$). When we further compress the V2X-ViT transmission to a level similar to UniV2X with Where2comm, there is a significant performance drop (from 0.287 to 0.071 at AMOTA). These outcomes underscore the capability of our infrastructure agent queries and agent fusion module in enhancing agent perception ability with light transmission cost.
Online Mapping Results.
We implement No Fusion, Early Fusion, and CoBEVT(Xu et al. 2022a) for online mapping on DAIR-V2X. All methods are re-implemented using BEVFormer (Li et al. 2022b). The mapping performance is reported with Segmentation Intersection over Union (IoU) (%) as the evaluation metric in Table 4. UniV2X demonstrates notable improvements in lane perception and crossing perception compared No Fusion, Early Fusion and CoBEVT, respectively. Moreover, compared with Early Fusion and CoBEVT, UniV2X requires less than 1/10th of the transmission cost. These results indicate that infrastructure lane queries and cross-view lane fusion are effective in enhancing online mapping ability.
@@
CoBEVT 15.6 16.4 2.56×$\times 10^{6}$
UniV2X (Ours) 17.8 (+11.4) 19.8 (+17.1) 1.47×$\times 10^{5}$
Occupancy Prediction Results.
Concerning the evaluation of occupancy prediction, as depicted in Table 5, UniV2X exhibits notably superior performance compared to No Fusion in both near and far regions. Particularly, UniV2X achieves +5.7 and +13.4 improvement in IoU-n (%) and IoU-f (%) respectively. Here, ”IoU-n” and ”IoU-f” denote evaluation ranges of 30×30m and 50×50m, respectively. These results underscore the effectiveness of our sparse-dense hybrid transmitted data in significantly enhancing occupancy prediction.
@@
Ablation Study on Reliability
When assessing UniV2X, our initial step involves randomly discarding 10%, 30%, 50%, 70%, and 100% of infrastructure agent queries during transmission to simulate data corruption. Following this, the retained queries are utilized for cross-view query transmission and interaction, and the performance of agent perception, encompassing object detection and tracking, is evaluated accordingly. The evaluation results depicted in Figure 4 unveil a gradual degradation in agent perception as the data corruption ratio increases. When the data corruption ratio reaches 100%, meaning only ego-vehicle sensor data can be used, the performance of UniV2X becomes comparable to that of the No Fusion model. This decline in performance is anticipated, as data corruption diminishes the complementary information crucial for ego-vehicle autonomous driving. Moreover, even in the absence of certain data due to corruption, UniV2X can maintain a basic level of performance comparable to the No Fusion model. This underscores the reliability of our transmission and cross-view data fusion mechanism.
Figure 4: Reliability on Data Corruption.
Conclusion
Limitations and Future Work. The framework involves multiple modules and different agent perspectives, resulting in a high degree of complexity. As a result, several interaction fusion modules within the framework remain in preliminary stages. Further refinement is essential for optimizing the internal design of the subsequent framework. In this work, we only consider open-loop evaluation for end-to-end autonomous driving. We will conduct more closed-loop experiments to evaluate our UniV2X.
Acknowledgment
This work is supported by the Wuxi Research Institute of Applied Technologies at Tsinghua University under Grant No. 20242001120, as well as the General Research Fund of Hong Kong under Grants No. 17200622 and 17209324.
Appendix
@@ Acknowledgment
1
Appendix A Related Works
Cooperative Autonomous Driving.
Leveraging V2X communication for cooperative autonomous driving has garnered significant attention. Works like (Wang et al. 2020; Li et al. 2021; Chen et al. 2019; Yang et al. 2023a; Hu et al. 2023a; Lin et al. 2024; Teufel et al. 2023; Yang et al. 2024) emphasize transmitting Bird’s Eye View (BEV) features extracted from point clouds or images to improve 3D object detection or tracking performance. (Xu et al. 2022a) employs sparse transformers to enhance segmentation performance, while(Ruan et al. 2023) focuses on utilizing sequential trajectories for motion forecasting. Addressing the latency issue in cooperation detection, (Lei et al. 2022; Yu et al. 2023b; Sizhe et al. 2023; Yang et al. 2023c) utilize historical frames or features. Additionally, (Hu et al. 2022; Fan et al. 2024; Chen, Shi, and Jia 2023; Zhong et al. 2024) transmit instance-level features or queries to reduce transmission costs and alleviate communication challenges. Datasets for cooperative object perception are provided by (Yu et al. 2022, 2023c; Xu et al. 2022b; Li et al. 2022a; Mao et al. 2022; Creß et al. 2022; Xu et al. 2022c).  (Yu et al. 2023c) releases a real-world trajectory dataset generated from infrastructure and vehicle sensor data. While most of these works concentrate on single tasks, a few also focus on end-to-end output planning or control tasks. For instance, (Valiente et al. 2019) employs multi-vehicle images with a simple convolutional network to generate control outputs. Similarly, (Cui et al. 2022) adopts a simple MLP to learn planning outputs with BEV features from multi-vehicle point clouds. However, these solutions lack explicit interpretability modules. In this paper, we present UniV2X, a comprehensive framework that integrates all essential modules within a single model, utilizing end-to-end learning for optimizing final planning improvement as well as intermediate outputs. Recently, several studies (Liu et al. 2024; Chen et al. 2024) have developed benchmarks to evaluate cooperative autonomous driving using closed-loop methodologies. In this work, we focus solely on open-loop evaluation for end-to-end autonomous driving. We plan to conduct closed-loop experiments to assess UniV2X in future work.
End-to-End Autonomous Driving.
End-to-end autonomous driving involves the extraction of planning output directly from raw data in a differentiable and learnable manner. Pioneering work, such as (Bojarski et al. 2016), employs CNNs to generate control outputs from point cloud data. Others, exemplified by (Sadat et al. 2020; Zeng et al. 2020), utilize point clouds and High-definition Maps as inputs. Concurrently, works like (Shao et al. 2023) leverage multi-modal sensor data as input to generate object density maps for visualization. For acquiring driving skills, (Codevilla et al. 2018; Prakash, Chitta, and Geiger 2021; Wu et al. 2022) employ imitation learning (IL) to learn from expert demonstrations in an open-loop manner. In contrast, (Liang et al. 2018; Kendall et al. 2019; Jia et al. 2023) utilize reinforcement learning (RL) to iteratively learn driving skills by interacting with the environment in a closed-loop fashion. Among related works, UniAD (Hu et al. 2023b) is the first to use queries to connect all essential tasks such as perception, mapping, prediction, and planning. Through unifying these tasks into a single network and employing imitation learning, UniAD achieves remarkable performance on the nuScenes dataset (Caesar et al. 2020). However, UniAD solely considers single-vehicle sensor data, while our approach, UniV2X, leverages sensor data from diverse views to enhance the overall driving ability.
Appendix B UniV2X Training
In our work, we train UniV2X with a four-stage training strategy to ensure comprehensive and stable training.
In the first stage, we pre-train the infrastructure and ego-vehicle systems separately. Specifically, we pre-train the infrastructure system, which includes tasks such as tracking, online mapping, and occupancy prediction, using annotations from the infrastructure view as ground truth. Similarly, we pre-train the ego-vehicle system, which involves tasks like tracking, online mapping, motion prediction, occupancy prediction, and planning, using annotations from the ego-vehicle view as ground truth.
In the second stage, we train the AgentFusion and LaneFusion modules. This stage encompasses tasks such as tracking and online mapping, utilizing cooperative-view annotations as ground truth, with the following combined loss:
$L=L_{\text{track}}+L_{\text{map}}.$ (5)
In the third stage, we train all fusion modules and all modules within the ego-vehicle system, covering all tasks. Cooperative-view annotations are used as ground truth for this stage, with the following combined loss:
$L=L_{\text{track}}+L_{\text{map}}+L_{\text{motion}}+L_{\text{occ}}+L_{\text{plan}}.$ (6)
Furthermore, we use self-supervised learning to train QueryFlowNet and OccFlowNet. This involves constructing infrastructure frame pairs and employing a similarity loss for training, similar to the strategy described in  (Yu et al. 2023b). This training requires no extra annotations.
Note that there is data distribution shift between the generated data from the clips selected from sequences in training stage and generated data from complete sequences in inference stage, especially for the track queries involving complex sequential queries correlation and update in TrackFormer. We found that if we train the agent fusion with the infrastructure agent queries generated from training clips, there will be unstable performance in inference stage. Thus we save the infrastructure agent queries of training part as the inference way, and train the agent fusion with the saved infrastructure agent queries. We are also exploring more friendly way to train the agent fusion module.
Appendix C Evaluation Metric Details
Object Detection.
We evaluate the detection performance using the mean Average Precision (mAP) metric, which is commonly used in previous works such as (Yu et al. 2022; Caesar et al. 2020). The AP is calculated based on the 11-points interpolated precision-recall curve and is defined as follows:
Multi-object Tracking.
AMOTA and AMOTP, 3D tracking evaluation metrics (Caesar et al. 2020), are used to measure the performance of multi-object tracking.
Online Mapping.
Occupancy Prediction.
In line with UniAD (Hu et al. 2023b), we assess the accuracy of predicted occupancy using a methodology similar to that outlined in (Hu et al. 2021; Zhang et al. 2022). Specifically, the Intersection over Union (IoU) metric quantifies categorical segmentations of the entire scene in an instance-agnostic manner. This metric is computed within Bird’s Eye View (BEV) ranges, with evaluations conducted for both near (30m×$\times$30m) and far (50m×$\times$50m) distances.
Planning.
We utilize the L2 error, collision rate and off-road rate metrics at different timestamps to assess the planning performance. Compared to existing open-loop evaluations, we have included the off-road rate as an additional metric. This metric is vital for evaluating planning performance, particularly in assessing how well the driving area is maintained and lane violations are minimized.
Transmission Cost.
Appendix D Ablation Study for Fusion Modules
In this part, we investigate the contributions of each fusion module within UniV2X. In evaluating the trained UniV2X, we only fuse agent queries, lane queries, and occupancy probability maps, respectively. We assess each individual fusion module and present the results in Table 6, 7 and 8. These tables illustrate that agent fusion, lane fusion, and occupancy fusion each contribute uniquely to agent perception, online mapping, and occupancy prediction, respectively. Additionally, all fusion modules collectively enhance the overall planning performance.
Table 6: Evaluation Results of UniV2X with No Fusion and Agent Fusion-Only.
Method Agent Perception Planning
mAP (%) $\uparrow$ AMOTA (%) $\uparrow$ Col. Rate (%, Avg.) $\downarrow$ Off-Road Rate (%, Avg.) $\downarrow$
No Fusion 16.5 16.3 0.89 1.08
Agent Fusion-Only 29.5 23.9 0.59 0.79
Table 7: Evaluation Results of UniV2X with No Fusion and Lane Fusion-Only.
Method Online Mapping Planning
IoU-Lane (%) $\uparrow$ IoU-CrossWalk (%) $\uparrow$ Col. Rate (%, Avg.) $\downarrow$ Off-Road Rate (%, Avg.) $\downarrow$
No Fusion 6.4 2.7 0.89 1.08
Lane Fusion-Only 17.8 19.8 0.59 0.89
Table 8: Evaluation Results of UniV2X with No Fusion and Occ Fusion-Only.
Method Occupancy Prediction Planning
IoU-n (%) $\uparrow$ IoU-f (%) $\uparrow$ Col. Rate (%, Avg.) $\downarrow$ Off-Road Rate (%, Avg.) $\downarrow$
No Fusion 16.3 13.1 0.89 1.08
Occ. Fusion-Only 20.3 23.5 0.39 0.98
Appendix E Ablation Study for Flow Prediction
Experiment Settings.